OpenStudy (anonymous):

If volume of a sphere increases by 72.8% , how much percent does it's Surface Area increases ?? If volume of a sphere increases by 72.8% , how much percent does it's Surface Area increases ?? @Mathematics

OpenStudy (anonymous):

well, the volume of a sphere is \[V=\frac{4}{3}\pi r^3\] the suface area is equal to \[SA=4\pi r^2\] start by determining how much r increases if you increase so\[1.728\frac{4}{3}\pi r^3=\frac{4}{3}\pi R^3\] and solve for R in terms of r once you determine R in terms of r, substitute R into the SA equation

OpenStudy (dumbcow):

first thing is you need to find how much the radius changed yeah thats exactly what i would say

OpenStudy (dumbcow):

for SA \[(1+p)4 \pi r^{2} = 4 \pi R^{2}\] you are solving for p, the percent increase in surface area